# Generalizing the SINDy approach with nested neural networks

### v1

@@ 5.2.1. Case 1: A trigonometric function involving composition using the PR block
=$\mathcal{F}=\{\ x\mapsto x,\ x\mapsto x^{2},\ x\mapsto x^{3},\ \mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits,\ x\mapsto\log(|x|),\ x\mapsto\sqrt{|x|}\ \}.$
@@
  • Optimizer: we use the Adam optimizer, with learning rate set to $0.01$ and multiplied by $0.999$ every $10$ epochs. We add noise to the gradient when the loss is stuck. We train the model for $100$ epochs.
@@
  • Sparsity: the Lasso parameter is oscillating around $10^{-4}$, and pruning parameters are set to $\text{MSE}_{\text{prune}}=0.01$, $\varepsilon_{\text{prune}}=0.01$, and $n_{\text{prune}}=2$.
The results for 10 and 20 epochs are displayed in Figure 8, where we observe that the model has not yet converged towards a satisfactory approximation. The model converges to the right formula after 100 epochs, as depicted on Figure 9.
@@
Figure 8. Case 1: training of Nested SINDy to recover equation (5.6). In the left figures, we compare $f$ (blue curve) and $f_{\theta}$ (orange curve), while in the right figures, we compare the trajectories of dynamic systems (5.5) (red curves) and (5.6) (blue curves) for three different initial conditions. After 20 epochs, the model still has not converged towards a satisfactory approximation.
@@
Figure 9. Case 1: training of Neural Nested SINDy to recover equation (5.6). In the left figures, we compare $f$ (blue curve) and $f_{\theta}$ (orange curve), while in the right figures, we compare the trajectories of dynamic systems (5.5) (red curves) and (5.6) (blue curves) for three different initial conditions. We observe that the model converges to the right formula after 100 epochs.
This example showcases the PR model’s ability to accurately learn the velocity of an ODE from trajectories. The next step is to conduct tests by increasing the dictionary size and using reduced data (fewer trajectories), sparser data (lower sampling frequency in the data), and noisy data.
@@ 5.2.2. Case 2: The Gompertz model
We train the model for $2000$ epochs. Results are summarized in Figure 11 and Table 2. After $1900$ epochs, the model converges to the right formula. During the training, the MSE between $f$ and $f_{\theta}$ is not always decreasing, but parameters are (slowly) pruned, and finally the model converges.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
As expected (see Section 5.1.3), PRP models take more effort to converge towards the expected function. In this case we succeed to recover the true formula of growth speed, but we considered a compact architecture with few dictionary functions, and $20$ times more epochs were required to converge than in the case of the PR model used in the previous case. As a consequence, a major way of improving the PRP models lies in improving the optimization process to make the convergence of such models easier.
@@ 6. Conclusion
In this study, we explored the capabilities and limitations of the Nested SINDy approach in discovering symbolic representations of dynamical systems from data. Our investigation covered a spectrum of cases, including the identification of “simple” functions, as well as the discovery of ordinary differential equations (ODEs) from trajectory data. The versatility of the Nested SINDy framework was demonstrated through various examples, showcasing its ability to approximate complex functions and discover ODEs underlying data-generating processes.
@@
Our results confirm that Nested SINDy can recover accurate symbolic expressions for a range of problems, from simple trigonometric functions to the more complex Gompertz model of tumor growth. However, the complexity of the optimization landscape associated with the Nested SINDy approach highlights the need for careful selection of hyperparameters and initialization strategies to ensure convergence to meaningful solutions.
Future work could focus on several areas to enhance the Nested SINDy framework. First, exploring alternative optimization algorithms specifically designed for the unique challenges of nested symbolic regression could improve the efficiency and reliability of the method. Additionally, incorporating mechanisms for automatic selection of the dictionary of basis functions based on preliminary data analysis might streamline the model development process and improve the model’s adaptability to different types of dynamical systems. Furthermore, one could envision integrating a supervised learning component, as in [45], in which a model is pre-trained to map the relationship between datasets and SINDy-like sparse patterns encoding analytical expressions. This effectively enables the automatic formulation of high quality initial solutions that can then be refined on a case-by-case basis. Finally, extending the Nested SINDy approach to handle partial differential equations (PDEs) could open new avenues for discovering the underlying physics of spatially extended systems from data.

### v2

@@ 5.2.1. Case 1: A trigonometric function involving composition using the PR block
=$\mathcal{F}=\{\ x\mapsto x,\ x\mapsto x^{2},\ x\mapsto x^{3},\ x\mapsto\mathop{\vphantom{\operator@font sin}}\!\mathop{\smash{\operator@font sin}\vphantom{s}}\nolimits(x),\ x\mapsto\log(|x|),\ x\mapsto\sqrt{|x|}\ \}.$
@@
  • Optimizer: we use the Adam optimizer, with learning rate set to $0.01$ and multiplied by $0.999$ every $10$ epochs. We add noise to the gradient when the loss do not decrease since $100$ epochs, as explained in (4.2) with $\alpha=16$.
@@
  • Sparsity: the Lasso parameter is oscillating around $\lambda_{0}=10^{-4}$ (see (4.1) for details), and pruning parameters are set to $\text{MSE}_{\text{prune}}=0.01$, $\varepsilon_{\text{prune}}=0.01$, and $n_{\text{prune}}=2$.
We train the model for $100$ epochs. The results for 10 and 20 epochs are displayed in Figure 8, where we observe that the model has not yet found a satisfactory approximation. The model retrieves the right formula after 100 epochs, as depicted on Figure 9 and in Table 2.
@@
Figure 8. Case 1 from Section 5.2.1: training of Nested SINDy to recover equation (5.7). In the left figures, we compare $f$ (blue curve) and $f_{\theta}$ (orange curve), while in the right figures, we compare the trajectories of dynamic systems (5.6) (red curves) and (5.7) (blue curves) for three different initial conditions. After 20 epochs, the model is still far from the true formula.
@@
Figure 9. Case 1 from Section 5.2.1: training of Neural Nested SINDy to recover equation (5.7). In the left figures, we compare $f$ (blue curve) and $f_{\theta}$ (orange curve), while in the right figures, we compare the trajectories of dynamic systems (5.6) (red curves) and (5.7) (blue curves) for three different initial conditions. We observe that the model approximates well the right formula after 100 epochs.
This example showcases the PR model’s ability to accurately learn the velocity of an ODE from trajectories. To further validate this ability, future research directions include conducting tests by increasing the dictionary size and using reduced data (fewer trajectories), sparser data (lower sampling frequency in the data), and noisy data.
Table 2. Case 1 from Section 5.2.1: evolution of the training of Neural Nested SINDy to recover equation (5.7). The table contains the formulas of $f_{\theta}$ for epochs greater than $50$, the mean squared error between $f_{\theta}$ and $f$, and the number of pruned parameters in the PRP block, at epochs 10, 20, 50, 100.
Epoch $f_{\theta}$ MSE on $[-3,3]$ MSE on $[-5,5]$ pruned parameters
10 expression too long to display $3.67\times 10^{-1}$ $4.16\times 10^{-1}$ 12
20 expression too long to display $2.91\times 10^{-1}$ $4.16\times 10^{-1}$ 14
[…]
@@ 5.2.2. Case 2: The Gompertz model
We train the model for $2000$ epochs. Results are summarized in Figure 11 and Table 3. After $1900$ epochs, the model approximates up to an error of $10^{-2}$ the right formula. During the training, the MSE between $f$ and $f_{\theta}$ is not always decreasing, but parameters are (slowly) pruned, and finally the model manages to approximate the right formula.
@@
As expected (see Section 5.1.3), PRP models take more effort to retrieve the expected function. In this case we succeed to approximate the true formula of growth speed, but we considered a compact architecture with few dictionary functions, and $20$ times more epochs than in the case of the PR model used in the previous case. As a consequence, a major way of improving the PRP models lies in improving the optimization process.
@@ 6. Conclusion
In this study, we explored the capabilities and limitations of the Nested SINDy approach in discovering symbolic representations of dynamical systems from data. Our investigation covered a spectrum of cases, including the identification of “simple” functions, as well as the discovery of ordinary differential equations (ODEs) from trajectory data. The versatility of the Nested SINDy framework was demonstrated through various examples, showcasing its ability to approximate complex functions and discover the ODEs underlying data-generating processes.
@@
Our results confirm that Nested SINDy can recover accurate symbolic expressions for a range of problems, from simple trigonometric functions to the more complex Gompertz model of tumor growth. However, the complexity of the optimization landscape associated with the Nested SINDy approach highlights the need for careful selection of hyperparameters and initialization strategies to ensure approximation of meaningful solutions.
Future work could focus on several areas to enhance the Nested SINDy framework. First, exploring alternative optimization algorithms specifically designed for the unique challenges of nested symbolic regression could improve the efficiency and reliability of the method. Additionally, incorporating mechanisms for automatic selection of the dictionary of basis functions based on preliminary data analysis might streamline the model development process and improve the model’s adaptability to different types of dynamical systems. Furthermore, one could envision integrating a supervised learning component, as in [45], in which a model is pre-trained to map the relationship between datasets and SINDy-like sparse patterns encoding analytical expressions. This effectively enables the automatic formulation of high-quality initial solutions that can then be refined on a case-by-case basis. Finally, extending the Nested SINDy approach to handle partial differential equations (PDEs) could open new avenues for discovering the underlying physics of spatially extended systems from data.